# Niobium coaxial cavities with internal quality factors exceeding 1.5 billion for circuit quantum electrodynamics

### v1

@@ Abstract
Group-V materials such as niobium and tantalum have become popular choices for extending the performance of circuit quantum electrodynamics (cQED) platforms allowing for quantum processors and memories with reduced error rates and more modes. The complex surface chemistry of niobium however makes identifying the main modes of decoherence difficult at millikelvin temperatures and single-photon powers. We use niobium coaxial quarter-wave cavities to study the impact of etch chemistry, prolonged atmospheric exposure, and the significance of cavity conditions prior to and during cooldown, in particular niobium hydride evolution, on single-photon coherence. We demonstrate cavities with quality factors of in the single-photon regime, a fold improvement over aluminum cavities of the same geometry. We rigorously quantify the sensitivity of our fabrication process to various loss mechanisms and demonstrate a reduction in the two-level system (TLS) loss tangent and a improvement in the residual resistivity over traditional BCP etching techniques. Finally, we demonstrate transmon integration and coherent cavity control while maintaining a cavity coherence of . The accessibility of our method, which can easily be replicated in academic-lab settings, and the demonstration of its performance mark an advancement in 3D cQED.
@@ I Introduction
An avenue for increasing the cavity coherence times of 3D cQED systems is to make the cavities out of bulk niobium. Elliptical niobium cavities used for accelerator applications have recently demonstrated single-photon lifetimes of for a cavity, with loaded lifetimes of approximately for a cavity after reducing TLS loss at the surface Romanenko et al. (2020a); Romanenko and Schuster (2017), showing the potential for such cavities in creating a next-generation of quantum memories. Recent results have also demonstrated qubit integration into elliptical cavities without severe degradation to overall cavity performance Milul et al. (2023).
@@ III Temperature Dependence of
To understand the effect that etch chemistry has on performance, two buffered chemical polish (BCP) recipes were used; one with the traditional 1:1:2 ratio of HF:HNO3:H3PO4, and one with 1:1:2 HF:HNO3:H2O. The results, shown in Fig. 1 (a), show cavities etched using H2O based BCP had improvement in single-photon coherence, and lower two-level system loss and residual resistivity when compared with H3PO4. This was even true for cavities that had previously been etched with H3PO4 BCP before being re-etched using the H2O buffered mixture, showing that performance was not intrinsic to a cavity, but a function of etch chemistry. This difference in performance and evidence of the underlying mechanism is discussed in Sec. IV.
@@
By studying the temperature or power dependence of the internal quality we can extract the two-level system loss tangent (), and the residual resistivity () simultaneously. The temperature dependence of was used due to the linearity of the cavity at drive strengths of photons. For these measurements the mixing chamber plate was stabilized to a temperature accuracy of prior to conducting time-domain spectroscopy.
@@
The results, shown in Fig. 2 (b), show a monotonic decrease in quality factor after only several temperature cycles. This is congruent with the lowering of due to NbHx formation at the surface due to the presence of hydrogen interstitials introduced during the etching process Knobloch (2003). Single-photon quality factor decreased by 36 over the same range. To determine the cause, time-of-flight secondary ion mass spectroscopy (ToF-SIMS) of the cavity’s center pin, taken post-cooldown and shown in Fig. 2 (c), shows a large hydrogen background extending deep into the bulk. By comparison, oxygen content becomes vanishingly small beyond . In SRF cavities this effect has been well characterized, with experiment suggesting cooldown rates in excess of 1K/min needed to avoid NbHx formation Knobloch (2003); Antoine and Berry (2003). By comparison, all cavities in this paper, except for Nb2-1, were initially cooled at rates of between without significant impact on performance, even after subsequent cooldowns, with temperature-dependent performance being consistent with elliptical SRF Nb cavities cooled in dilution refrigerators Romanenko et al. (2020a).
With NbHx evolution explored, a third cavity was cooled to better investigate the evolution of the oxide and subsequent two-level system (TLS) loss mechanisms during and after the H2O etching process. Niobium is known to host TLS in its oxide layer, which is believed to be caused by the presence of NbOx states within a disordered thick Nb2O5 surface oxide Romanenko and Schuster (2017); Premkumar et al. (2021). To quantify this the extracted temperature-dependent quality factors were fit to the following TLS temperature model Gao et al. (2008); Lei et al. (2020); Romanenko et al. (2020a):
Where , with and being the surface to volume participation ratio of the electric field and surface dielectric thickness respectively. , where is the magnetic surface to volume participation ratio, and is the residual surface resistivity. The loss tangent product, for the H2O buffered process ranged from when hermetically sealed within 30 minutes of etch quenching, which is as good or better than state of the art EP etched elliptical cavities prior to additional oxide removal Romanenko et al. (2020b), and better than H3PO4 buffered etching.
@@
To understand the relative concentrations of compounds at the surface, the XPS spectra, shown in Fig. 3 (a), was fit for both freshly etched and unsealed cavities that had been exposed to the atmosphere for prolonged periods of time. The fit spectra reveal four oxidation states and the presence of NbHx, particularly in the underlying Nb-metal surface, which are absent in the Nb thin-film shown in Fig. S4 (c), that was etched using fluorine reactive ion etching instead of using an acid based etch. The relative ratios of the Nb metal phase to the Nb2O5, NbO, NbO2, and disordered NbOx suggest a structural change in the oxide. Using techniques outlined in, Appendix E, these ratios were also used to determine oxide thickness Premkumar et al. (2021); Soda et al. (2013). From the fits, the oxide layer grew by 38, from to . These oxide thicknesses are comparable to previously reported values in un-annealed electropolished (EP) cavities Romanenko and Schuster (2017); Premkumar et al. (2021). Given the linear dependence of the TLS loss-tangent with oxide thickness, this increase does not fully account for the 75 increase in measured loss tangent product in Fig. 3 (a). This implies that oxide morphology and composition play a comparable role in the measured loss.
@@ V Discussion
Surface analysis using XPS and ToF-SIMS shows both the significant introduction of hydrogen to depths of and the strong dependence of exposure time to atmosphere, increased oxide thickness, and a subsequent increase in the TLS loss of the resonator. Rapid sealing and evacuation () seems to keep the TLS loss product at values comparable to electropolished Nb elliptical cavities before any additional mid-T annealing to diffuse the oxide Romanenko et al. (2020b); Romanenko and Schuster (2017). This is notable since the geometric participation is greater in the coaxial design than the mode of an elliptical cavity, with a filling fraction that is also greater than that reported for TESLA designs Romanenko et al. (2020b). This means that, when scaled, the are within a factor of 3 of the state-of-art Nb elliptical cavities of similar frequencies, and effective TLS loss-tangents are between lower than normal EP treated cavities with no additional oxide annealing Romanenko et al. (2020b). XPS analysis in Fig. 3 (a) seems to hint at the source of this discrepancy, with the H2O BCP etched cavities showing a thin oxide () that is largely composed of Nb2O5. The use of HNO3, which is a strong oxidant, may preferentially produce pentoxide over other suboxides, producing a lower-loss dielectric surface.
Another insight was the effect that hydrogen interstitials and NbHx formation play on single-photon performance, with NbHx formation not leading to a significant decrease in single-photon at cooldown rates between . While annealing would improve , the above data shows a peak in at temperatures consistent with cavities that have undergone annealing, suggesting that the cavity is not noticeably affected after a single cooldown cycle Romanenko et al. (2020b).
@@ Acknowledgements.
The authors would like to acknowledge Alexander Filatov for his assistance in XPS analysis and Aaron Chou for his programmatic support of this project. ToF-SIMS measurements were done with assistance from Eurofins EAG Laboratories. We would also like to acknowledge Justin Jureller and Luigi Mazzenga for technical assistance throughout. This work was supported by the Samsung Advanced Institute of Technology Global Research Partnership. This work is also supported by the U.S. Department of Energy Office of Science National Quantum Information Science Research Centers as part of the Q-NEXT center. Material fabrication and characterization was partially supported by the University of Chicago Materials Research Science and Engineering Center, which is funded by the National Science Foundation under award number DMR-2011854. Transmon qubit devices were fabricated in the Pritzker Nanofabrication Facility at the University of Chicago, which receives support from Soft and Hybrid Nanotechnology Experimental (SHyNE) Resource (NSF ECCS-2025633), a node of the National Science Foundation’s National Nanotechnology Coordinated Infrastructure. This document was prepared using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Research Alliance, LLC (FRA), acting under Contract No. DE-AC02-07CH11359.Appendix A Cavity Manufacturing and Etching
The cavity was machined out of a piece of pure niobium metal with a residual resistivity ratio (RRR) of which conformed to ASTM B393 R04220 standards and was purchased from Ningxia Nonferrous Metals. For more even material removal during etching, sharp edges and corners of the co-axial design were radiused for smooth transitions. The cavity was initially degreased using an Alconox and DI-water mixture at temperatures under sonication. Following this, the cavity was thoroughly rinsed in DI-water and dried using an isopropyl alcohol (IPA) solvent exchange and dry nitrogen. A toluene-acetone-methanol-IPA cleaning was done using semiconductor (semi) grade solvents under sonication for at each step, before being dried under high-purity nitrogen that had been passed through a filter before being bagged in a ISO100 clean-bag. All cavities were stored in a desiccated dry box after cleaning and prior to etching.
@@
The etch recipe used in this paper diverges from the traditional BCP procedure, using a 1:1:2 HF:HNO3:H2O etch chemistry. The smaller surface area of our cavity, as listed in the Fig. S1 (a) table, means that our dissipated power is for even our highest etching rates. This, combined with the smaller volume of etchant required, means that an unbuffered etch can be used with minimal degradation in etch quality while keeping the reaction kinetics in a safe and manageable regime. In addition, the closed nature of the cavity geometry, with poor acid circulation, means that higher viscosity BCP mixtures can greatly affect the etch performance at the surface.
The main etching steps of the cavity are depicted in Fig. S1 Cav etch. A custom Teflon PTFE holder was built to facilitate rapid and safe transfer of the cavity between etching step (depicted in Fig. S1). PCTFE/PVDF screws were used to attach the cavity to the holder. The main etching was done in a PTFE beaker. The beaker was placed inside of a water bath, and surrounded by a copper coil attached to a thermoelectric cooler which circulates a water/glycol mixture through the coil. The whole water bath was set on a magnetic stirring plate to circulate the primary acid volume. A pure niobium test strip is placed into the etch container alongside the cavity. The strip is periodically removed, rinsed, and thickness measured using a micrometer to check etch-rate. Fig. S1 (a) provides a diagram of the aforementioned etch setup. To circulate the acid inside the cavity volume, the etchant is manually exchanged using a clean polypropylene pipette at min intervals. Other pertinent etching parameters are listed in the tables of Fig. S1 (ac).
@@
Following drying the cavity is sealed. An indium ring is sandwiched between the top of the cavity and the cap. A 3mm stainless steel tube is brazed into the copper cap. A Swagelok VCR fitting is brazed at the other end. A second VCR fitting is attached to a copper pinch-off tube, allowing for easy replacement and reuse of the cap. A Corning-Gilbert GPO (SMP) hermetic connector (Part 0119-783-1) is sealed into the top of the cap via a second indium o-ring, and allows for microwave feedthrough. An antenna is made on the vacuum-side of the feedthrough to couple RF power into the cavity. Because of the evanescent decay of the field energy, the coupling quality factor is exponentially sensitive to this antenna distance from the cavity center pin. The cavity is left under vacuum for before being sealed via a hydraulic pinch-off tool and placed into the fridge. A diagram of the sealed cavity with pinch-off tube is shown in Fig. S2 (b).
@@ Appendix D Surface Characterization
The compositions of both surface oxides and the underlying niobium metal were determined using X-ray photoemission spectroscopy (XPS). Analysis was performed on a Kratos Axis Nova spectrometer using a monochromatic Al K source (1486.6 eV). Nb 3d, C 1s, and O 1s high-resolution spectra were collected using an analysis area of mm and 20 eV pass energy with the step size of 100 meV. The survey spectra were collected using 100 eV pass energy. Charge neutralization was performed using a co-axial, low energy ( 0.1 eV) electron flood source to avoid shifts in the recovered binding energy. The C 1s peak of adventitious carbon was set at 284.8 eV to compensate for any remaining charge-induced shifts. Deconvolution of the high-resolution XPS spectra was performed in CasaXPS software using Lorentzian asymmetric curves for the Nb metal phase, symmetric Gaussian-Lorentzian curves for all other elements’ fitting, and a Shirley fitting for the background. The Nb 3d region consists of the two 3d5/2 and 3d3/2 spin-orbit split components. The peak area ratio of 3d5/2 to 3d3/2 was fixed to 3:2. The Nb 3d region was fit using six doublet components (Nb metal, Nb2O5, NbO2, NbO, NbOx, NbHx) of 3d5/2 and 3d3/2 components for each sample. The energy splitting of each doublet component is 2.7 eV.

### v2

@@ Abstract
Group-V materials such as niobium and tantalum have become popular choices for extending the performance of circuit quantum electrodynamics (cQED) platforms allowing for quantum processors and memories with reduced error rates and more modes. The complex surface chemistry of niobium however makes identifying the main modes of decoherence difficult at millikelvin temperatures and single-photon powers. We use niobium coaxial quarter-wave cavities to study the impact of etch chemistry, prolonged atmospheric exposure, and the significance of cavity conditions prior to and during cooldown, in particular niobium hydride evolution, on single-photon coherence. We demonstrate cavities with quality factors of in the single-photon regime, a fold improvement over aluminum cavities of the same geometry. We rigorously quantify the sensitivity of our fabrication process to various loss mechanisms and demonstrate a reduction in the two-level system (TLS) loss tangent and a improvement in the residual resistance over traditional BCP etching techniques. Finally, we demonstrate transmon integration and coherent cavity control while maintaining a cavity coherence of . The accessibility of our method, which can easily be replicated in academic-lab settings, and the demonstration of its performance mark an advancement in 3D cQED.
@@ I Introduction
An avenue for increasing the cavity coherence times of 3D cQED systems is to make the cavities out of bulk niobium. Elliptical niobium cavities used for accelerator applications have recently demonstrated single-photon lifetimes of for a cavity, with loaded lifetimes of approximately for a cavity after reducing TLS loss at the surface Romanenko et al. (2020); Romanenko and Schuster (2017), showing the potential for such cavities in creating a next-generation of quantum memories. Recent results have also demonstrated qubit integration into elliptical cavities without severe degradation to overall cavity performance Milul et al. (2023).
@@ III Temperature Dependence of
To understand the effect that etch chemistry has on performance, two buffered chemical polish (BCP) recipes were used; one with the traditional 1:1:2 ratio of HF:HNO3:H3PO4, and one with 1:1:2 HF:HNO3:H2O. The results, shown in Fig. 1 (a), show cavities etched using H2O based BCP had improvement in single-photon coherence, and lower two-level system loss and residual resistance when compared with H3PO4. This was even true for cavities that had previously been etched with H3PO4 BCP before being re-etched using the H2O buffered mixture, showing that performance was not intrinsic to a cavity, but a function of etch chemistry. This difference in performance and evidence of the underlying mechanism is discussed in Sec. IV.
@@
By studying the temperature or power dependence of the internal quality we can extract the two-level system loss tangent (), and the residual resistance () simultaneously. The temperature dependence of was used due to the linearity of the cavity at drive strengths of photons. For these measurements the mixing chamber plate was stabilized to a temperature accuracy of prior to conducting time-domain spectroscopy.
@@
The results, shown in Fig. 2 (b), show a monotonic decrease in quality factor after only several temperature cycles. This is congruent with the lowering of due to NbHx formation at the surface due to the presence of hydrogen interstitials introduced during the etching process Knobloch (2003). Single-photon quality factor decreased by 36 over the same range. To determine the cause, time-of-flight secondary ion mass spectroscopy (ToF-SIMS) of the Nb2-1 cavity’s center pin, taken post-cooldown and shown in Fig. 2 (c), shows a large hydrogen background extending deep into the bulk. By comparison, oxygen content becomes vanishingly small beyond . In SRF cavities this effect has been well characterized, with experiment suggesting cooldown rates in excess of 1K/min needed to avoid NbHx formation Knobloch (2003); Antoine and Berry (2003). By comparison, all cavities in this paper, except for Nb2-1, were initially cooled at rates of between without variation on performance, even after subsequent cooldowns. While the temperature dependence of is comparable to that of elliptical SRF cavities cooled to dilution temperatures Romanenko et al. (2020), higher overall residual resistance when compared to cavities that have undergone hydrogen removal still suggests some impact on performance Ciovati et al. (2010).
With NbHx evolution explored, a third cavity was cooled to better investigate the evolution of the oxide and subsequent two-level system (TLS) loss mechanisms during and after the H2O etching process. Niobium is known to host TLS in its oxide layer, which is believed to be caused by the presence of NbOx states within a disordered thick Nb2O5 surface oxide Romanenko and Schuster (2017); Premkumar et al. (2021); Bafia et al. (2024). To quantify this the extracted temperature-dependent quality factors were fit to the following TLS temperature model Gao et al. (2008); Lei et al. (2020); Romanenko et al. (2020):
Where , with and being the surface to volume participation ratio of the electric field and surface dielectric thickness respectively. , where is the magnetic surface to volume participation ratio, and is the residual surface resistivity. and were determined via Ansys HFSS to be and respectively. The loss tangent product, for the H2O buffered process ranged from when hermetically sealed within 30 minutes of etch quenching, which is as good or better than state of the art EP etched elliptical cavities prior to additional oxide removal Romanenko et al. (2020), and better than H3PO4 buffered etching.
@@
To understand the relative concentrations of compounds at the surface, the XPS spectra, shown in Fig. 3 (a), was fit for both freshly etched and unsealed cavities that had been exposed to the atmosphere for prolonged periods of time. The fit spectra reveal four oxidation states and the presence of NbHxKalboussi et al. (2025), particularly in the underlying Nb-metal surface, which are absent in the Nb thin-film shown in Fig. S4 (c), that was etched using fluorine reactive ion etching instead of using an acid based etch. The relative ratios of the Nb metal phase to the Nb2O5, NbO, NbO2, and Nb2O suggest a structural change in the oxide. Using techniques outlined in, Appendix E, these ratios were also used to determine oxide thickness Premkumar et al. (2021); Soda et al. (2013). From the fits, the oxide layer grew by 38, from to . These oxide thicknesses are comparable to previously reported values in unbaked electropolished (EP) cavities Romanenko and Schuster (2017); Premkumar et al. (2021). Given the linear dependence of the TLS loss-tangent with oxide thickness, this increase does not fully account for the 75 increase in measured loss tangent product in Fig. 3 (a). This implies that oxide morphology and composition play a comparable role in the measured loss.
@@ V Discussion
Surface analysis using XPS and ToF-SIMS shows both the significant introduction of hydrogen to depths of and the strong dependence of exposure time to atmosphere, increased oxide thickness, and a subsequent increase in the TLS loss of the resonator. Rapid sealing and evacuation () seems to keep the TLS loss product at values comparable to electropolished Nb elliptical cavities before any additional mid-T baking to diffuse the oxide Romanenko et al. (2020); Romanenko and Schuster (2017). This is notable since the geometric participation is greater in the coaxial design than the mode of an elliptical cavity, with a filling fraction that is also greater than that reported for TESLA designs Romanenko et al. (2020). This means that, when scaled, the are within a factor of 3 of the state-of-art Nb elliptical cavities of similar frequencies, and effective TLS loss-tangents are between lower than normal EP treated cavities with no additional oxide bakingRomanenko et al. (2020). This, combined with an increase in TLS loss that is greater than what may be expected from increased oxide thickness alone as shown in Fig. 3 (a), suggests that H2O BCP etched cavities have an oxide with fewer oxygen vacancies than other wet-oxidation techniques Bafia et al. (2024); Grundner and Halbritter (1980). A careful accounting of excess oxygen counts in ToF-SIMs may be one way to rectify this difference between EP and the H2O BCP presented hereBafia et al. (2024).
Another insight was the effect that hydrogen interstitials and NbHx formation play on single-photon performance, with NbHx formation not leading to a significant decrease in single-photon at cooldown rates between over successive cooldowns. Even when hydride formation is induced, decrease by only 36%. While baking would improve , the above data shows a peak in at temperatures consistent with cavities that have undergone baking, suggesting that the cavity is not noticeably affected after a single cooldown cycle Romanenko et al. (2020).
@@ Acknowledgements.
The authors would like to acknowledge Alexander Filatov for his assistance in XPS analysis and Aaron Chou for his programmatic support of this project. ToF-SIMS measurements were done with assistance from Eurofins EAG Laboratories. We would also like to acknowledge Justin Jureller and Luigi Mazzenga for technical assistance throughout. This work was supported by the Samsung Advanced Institute of Technology Global Research Partnership. This work is also supported by the U.S. Department of Energy Office of Science National Quantum Information Science Research Centers as part of the Q-NEXT center. Material fabrication and characterization was partially supported by the University of Chicago Materials Research Science and Engineering Center, which is funded by the National Science Foundation under award number DMR-2011854. Transmon qubit devices were fabricated in the Pritzker Nanofabrication Facility at the University of Chicago, which receives support from Soft and Hybrid Nanotechnology Experimental (SHyNE) Resource (NSF ECCS-2025633), a node of the National Science Foundation’s National Nanotechnology Coordinated Infrastructure. This document was prepared using the resources of the Fermi National Accelerator Laboratory (Fermilab), a U.S. Department of Energy, Office of Science, Office of High Energy Physics HEP User Facility. Fermilab is managed by Fermi Research Alliance, LLC (FRA), acting under Contract No. DE-AC02-07CH11359.
Appendix A Cavity Manufacturing and Etching
The cavity was machined out of a piece of pure niobium metal with a residual resistance ratio (RRR) of which conformed to ASTM B393 R04220 standards and was purchased from Ningxia Nonferrous Metals. For more even material removal during etching, sharp edges and corners of the co-axial design were radiused for smooth transitions. The cavity was initially degreased using an Alconox© and DI-water mixture at temperatures under sonication. Following this, the cavity was thoroughly rinsed in DI-water and dried using an isopropyl alcohol (IPA) solvent exchange and dry nitrogen. A toluene-acetone-methanol-IPA cleaning was done using semiconductor (semi) grade solvents under sonication for at each step, before being dried under high-purity nitrogen that had been passed through a filter before being bagged in a ISO100 clean-bag. All cavities were stored in a desiccated dry box after cleaning and prior to etching.
| Etching Parameters | |
| Etch | 1:1:2 (HF:HNO3:H2O)† |
| Etch depth (m) | |
| Etched area () | 70.2 |
| Time (min) | |
| Temp () | |
| Etch vol. (mL) | |
| Stir-rate (min-1) | |
| Pipette Interval (min) | |
| Quenching and Rinse Parameters | |
| Water resistivity () | |
| Water TOC (PPB) | 10 |
| Quench water vol () | |
| Quench time, 1st step (min) | |
| Quench time, 2nd step (min) | |
| Temp () | |
| DI rinse time (min) | |
| Drying solvent | IPA (semi-grade) |
| Solvent exchange time (min) | |
| Drying gas | N2§ |
@@ Appendix A Cavity Manufacturing and Etching
The etch recipe used in this paper diverges from the traditional BCP procedure, using a 1:1:2 HF:HNO3:H2O etch chemistry. The smaller surface area of our cavity, as listed in Table 1, means that our dissipated power is for even our highest etching rates. This, combined with the smaller volume of etchant required, means that an unbuffered etch can be used with minimal degradation in etch quality while keeping the reaction kinetics in a safe and manageable regime. In addition, the closed nature of the cavity geometry, with poor acid circulation, means that higher viscosity BCP mixtures can greatly affect the etch performance at the surface.
The main etching steps of the cavity are depicted in Fig. S1 Cav etch. A custom Teflon© PTFE holder was built to facilitate rapid and safe transfer of the cavity between etching step (depicted in Fig. S1). PCTFE/PVDF screws were used to attach the cavity to the holder. The main etching was done in a PTFE beaker. The beaker was placed inside of a water bath, and surrounded by a copper coil attached to a thermoelectric cooler which circulates a water/glycol mixture through the coil. The whole water bath was set on a magnetic stirring plate to circulate the primary acid volume. A pure niobium test strip is placed into the etch container alongside the cavity. The strip is periodically removed, rinsed, and thickness measured using a micrometer to check etch-rate. Fig. S1 (a) provides a diagram of the aforementioned etch setup. To circulate the acid inside the cavity volume, the etchant is manually exchanged using a clean polypropylene pipette at min intervals. Other pertinent etching parameters are listed in the tables of Fig. S1 (ac).
@@
Following drying the cavity is sealed. An indium ring is sandwiched between the top of the cavity and the cap. A 3mm stainless steel tube is brazed into the copper cap. A Swagelok VCR© fitting is brazed at the other end. A second VCR fitting is attached to a copper pinch-off tube, allowing for easy replacement and reuse of the cap. A Corning-Gilbert GPO (SMP) hermetic connector (Part 0119-783-1) is sealed into the top of the cap via a second indium o-ring, and allows for microwave feedthrough. An antenna is made on the vacuum-side of the feedthrough to couple RF power into the cavity. Because of the evanescent decay of the field energy, the coupling quality factor is exponentially sensitive to this antenna distance from the cavity center pin. The cavity is left under vacuum for before being sealed via a hydraulic pinch-off tool and placed into the fridge. A diagram of the sealed cavity with pinch-off tube is shown in Fig. S2 (b).
@@ Appendix D Surface Characterization
The compositions of both surface oxides and the underlying niobium metal were determined using X-ray photoemission spectroscopy (XPS). Analysis was performed on a Kratos Axis Nova spectrometer using a monochromatic Al K source (1486.6 eV). Nb 3d, C 1s, and O 1s high-resolution spectra were collected using an analysis area of mm and 20 eV pass energy with the step size of 100 meV. The survey spectra were collected using 100 eV pass energy. Charge neutralization was performed using a co-axial, low energy ( 0.1 eV) electron flood source to avoid shifts in the recovered binding energy. The C 1s peak of adventitious carbon was set at 284.8 eV to compensate for any remaining charge-induced shifts. Deconvolution of the high-resolution XPS spectra was performed in CasaXPS software using Lorentzian asymmetric curves for the Nb metal phase, symmetric Gaussian-Lorentzian curves for all other elements’ fitting, and a Shirley fitting for the background. The Nb 3d region consists of the two 3d5/2 and 3d3/2 spin-orbit split components. The peak area ratio of 3d5/2 to 3d3/2 was fixed to 3:2. The Nb 3d region was fit using six doublet components (Nb metal, Nb2O5, NbO2, NbO, Nb2O, NbHx) of 3d5/2 and 3d3/2 components for each sample. The energy splitting of each doublet component is 2.7 eV.
| Depth | Sample | Parameters | Nb | Nb2O5 3d5/2 | NbO2 3d5/2 | NbO 3d5/2 | Nb2O 3d5/2 | NbHx 3d5/2 |
| Nb2-4 | Binding Energy (eV) | 202.3 | 207.95 | 206.1 | 204.1 | 203.2 | 203.89 | |
| +22 days | FWHM (eV) | 0.75 | 1.45 | 1.5 | 1.5 | 1.3 | 1.8 | |
| Atomic Comp. () | 38.1 | 52.59 | 2.2 | 0.98 | 0.54 | 5.47 | ||
| Nb2-4 | Binding Energy (eV) | 202.35 | 207.95 | 206.1 | 204.2 | 203.2 | 203.89 | |
| Surface | +30 min | FWHM (eV) | 0.75 | 1.45 | 1.50 | 1.5 | 1.3 | 1.8 |
| Atomic Comp. () | 60.55 | 23.78 | 4.37 | 7.05 | 0.62 | 2.58 | ||
| Nb | Binding Energy (eV) | 202.27 | 207.9 | 206.2 | 204.2 | 203.2 | — | |
| Film | FWHM (eV) | 0.75 | 1.4 | 1.50 | 1.40 | 1.40 | — | |
| Atomic Comp. () | 35.87 | 41.16 | 17.65 | 4.13 | 1.15 | — | ||
| Nb2-4 | Binding Energy (eV) | 202.4 | — | — | — | — | 203.90 | |
| +22 days | FWHM (eV) | 0.75 | — | — | — | — | 2.0 | |
| Atomic Comp () | 68.8 | — | — | — | — | 31.2 | ||
| Nb2-4 | Binding Energy (eV) | 202.4 | — | — | — | — | 203.9 | |
| +35nm depth | +30 min | FWHM (eV) | 0.75 | — | — | — | — | 2.0 |
| Atomic Comp. () | 68.1 | — | — | — | — | 31.9 | ||
| Nb | Binding Energy (eV) | 202.3 | — | — | — | — | — | |
| Film | FWHM (eV) | 0.6 | — | — | — | — | — | |
| Atomic Comp. () | 100 | — | — | — | — | — |